OpenStudy (anonymous):

what is the summation formula for x^1? if x^2,n(n+1)/6.

OpenStudy (anonymous):

ahh variables lemme see if i can figure it out

OpenStudy (anonymous):

nvm srry but ill give u a medal :)

OpenStudy (anonymous):

they wont let me

OpenStudy (anonymous):

this question could use some clarification what is \[x^1\] ? is it just \[x\]? wand what does if x^2, n(n+1)/6 mean?

OpenStudy (anonymous):

summation formula for X power 2 is n(n+1)/6 so what I want to know is,whether the summation formula X power 1 same with X power 2 or not

OpenStudy (anonymous):

ok first off i assume you mean \[\sum_{k=1}^{\infty}k^2\] right? and if so, it is not \[\frac{n(n+1)}{6}\] but rather it is \[\sum_{k=1}^{\infty}k^2=\frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

then,its same with X power 1?

OpenStudy (anonymous):

actually what i wrote was silly. i meant \[\sum_{k=1}^{n}k^2=\frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

sum from 1 to n

OpenStudy (anonymous):

for summing consecutive integers the formula is \[\sum_{k=1}^n k =\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

how to solve this,use summation are of definition of area to find area under y=1+3x on(-1,5)

OpenStudy (anonymous):

use which formula?

OpenStudy (anonymous):

are you looking for the limit of the reimann sum?

OpenStudy (anonymous):

area as the limit of sum

OpenStudy (anonymous):

what a pain.

OpenStudy (anonymous):

yeah,realy pain..

OpenStudy (anonymous):

first of all we can find the answer using geometry.

OpenStudy (anonymous):

the answer is 42 by plane geometry, so at least we know what we are looking for

OpenStudy (anonymous):

now we can do this the smart way i suppose. we want to take the limit of the sum. \[\lim_{n\rightarrow \infty} \sum_{k=1}^n f(x_k)\Delta x\]

OpenStudy (anonymous):

my \[\Delta x\] is 6/n

OpenStudy (anonymous):

we can at least break it apart into two pieces, since addition is commutative and also we need to write \[\Delta x=\frac{5+1}{n}=\frac{6}{n}\] \[\lim_{n\rightarrow \infty} \sum_{k=1}^n 1 \Delta x=\lim_{n\rightarrow \infty}\sum_{k=1}^n \frac{6}{n}\] \[=\frac{6}{n}\sum_{k=1}^n 1=6\]

OpenStudy (anonymous):

right we use \[\Delta x = \frac{6}{n}\]

OpenStudy (anonymous):

my fk\[\Delta x\] is 108k/n -12/n

OpenStudy (anonymous):

correct or not?

OpenStudy (anonymous):

are you breaking it in to two pieces, or just using \[1+3x\]?

OpenStudy (anonymous):

i would make my life easier and write the sum in two parts, since we have already computed the first part. it is 6

OpenStudy (anonymous):

then i would look at the function \[f(x)=x\] on (-1,5) and say \[x_0=-1, x_1=-1+\frac{6}{n}, x_2=-1+\frac{12}{n} ...x_k=-1+\frac{6k}{n}\]

OpenStudy (anonymous):

i realize we have to multiply by 3, but that comes right out front of the summation because \[\sum 3f(x)=3\sum f(x)\]

OpenStudy (anonymous):

so all that is left is to write \[\sum_{k=1}^n (-1+\frac{6k}{n})\times \frac{6}{n}\]

OpenStudy (anonymous):

oh times 3

OpenStudy (anonymous):

\[3\sum_{k=1}^n (-1+\frac{6k}{n})\times \frac{6}{n}\]

OpenStudy (anonymous):

why times 3?

OpenStudy (anonymous):

\[3\sum_{k=1}^n -\frac{6}{n}+\frac{36k}{n^2}\] \[-18+\frac{36}{n^2}\sum_{k=1}^n k\]

OpenStudy (anonymous):

oh because i want to make life easy

OpenStudy (anonymous):

we have \[f(x)=1+3x\] and so i am just saying we can sum up as \[\sum (1+3x) dx=\sum 1dx +3\sum x dx\]

OpenStudy (anonymous):

the first sum we already computed here \[=\frac{6}{n}\sum_{k=1}^n 1=6\]

OpenStudy (anonymous):

and it is just easier for me to work with \[f(x)=x\] than \[f(x)=3x\]although it really makes no difference the 3 comes right out front of the summation

OpenStudy (anonymous):

however i forgot the distributive law! it should be \[3\sum_{k=1}^n (-1+\frac{6k}{n})\times \frac{6}{n}\] \[=3\sum_{k=1}^n -\frac{6}{n}+\frac{36k}{n^2}\] \[=\frac{-18}{n}\sum_{k=1}^n 1+\frac{108}{n^2}\sum_{k=1}^n k\] \[=-18+\frac{108}{n^2}\sum_{k=1}^n k\] and only now do we use the summation formula \[\sum_{k=1}^n k =\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

\[\frac{108}{n^2}\frac{n(n+1)}{2}\] \[\frac{54(n^2+n)}{n^2}\] take the limit as n goes to infinity, get 54

OpenStudy (anonymous):

so we should get \[6-18+54\] as our answer, i hope it is right.. yup, \[6-18+54=42\]

OpenStudy (anonymous):

how to get 54?

OpenStudy (anonymous):

start with this line here \[=-18+\frac{108}{n^2}\sum_{k=1}^n k\] and ignoring the -18 we need to compute \[\frac{108}{n^2}\sum_{k=1}^n k\] now we use the summation formula (closed form of the sum) to get \[\frac{108}{n^2}\times \frac{n(n+1)}{2}\] clear so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we are done with summations completely. we have computed everything. all that is left to do is to take the limit as \[\Delta x \rightarrow 0\] i.e. as \[n\rightarrow \infty\]

OpenStudy (anonymous):

and if you look at this expression \[\frac{108}{n^2}\frac{n(n+1)}{2}=\frac{54(n^2+n)}{n^2}\] you want \[\lim_{n\rightarrow \infty}\frac{54(n^2+n)}{n^2}\]

OpenStudy (anonymous):

you have a rational function where the numerator and denominator both have degree 2, so you take the ratio of the leading coefficients and get that limit is 54 pretty much in your head

OpenStudy (anonymous):

i suppose you could write something out fancier like divide top and bottom by n^2 but that is a waste of time. you could even use l'hopital if you like but that is like killing a gnat with a 45

OpenStudy (anonymous):

if you want to understand further, since we have the answer, instead of breaking it in to two pieces you could keep the sum as \[\sum f(x_k)\frac{6}{n}=\sum 1+3(-1+\frac{6k}{n})\frac{6}{n}\] but the work will be identical

OpenStudy (anonymous):

i get like that

OpenStudy (anonymous):

oke2 i try again.

OpenStudy (anonymous):

then your next step will be to break it up in to pieces just like we did previous, and to compute just as before, so it will be repetition, but you will learn something by doing it

OpenStudy (anonymous):

how get the 6-18+54?

OpenStudy (anonymous):

until here,,i just get until 54

OpenStudy (anonymous):

where does 6-18+54 come from?

OpenStudy (anonymous):

6 from the first sum, -18 from the second part